The sum of two numbers is 208. One-third of the greater number equals the smaller number. What are the numbers?
Answer by glitzgirl_14(11)

For the Larger number I will use the variable 'L' and for the smaller I will use 's'.
L+s=208
(1/3)L=s
Put (1/3)L in for s in the first equation:
L+1/3L=208
1 1/3L=208
L=156
156+s=208
s=52